OpenStudy (anonymous):

1/R=1/R1+1/R2 for R1

OpenStudy (anonymous):

1/R1=1/R-1/R2 = R2-R/R.R2 R1=R*R2/R2-R

OpenStudy (anonymous):

Thank you for your assistance theyatin.